OpenStudy (anonymous):

solve elimination plz show work x+y=20 x-y=4

OpenStudy (anonymous):

plz i need help

OpenStudy (anonymous):

THERE ARE TWO STEPS ! FIRST STEP:- In x+y=20, their should be following cases: 1) x=0 and y=20 2) x=1 and y=19 3) x=2 and y=18 4) x=3 and y=17 5) x=4 and y=16 6) x=5 and y=15 7) x=6 and y=14 8) x=7 and y=13 9) x=8 and y=12 10) x=9 and y=11 11) x=10 and y=10 12) x=11 and y=9 13) x=12 and y=8 14) x=13 and y=7 15) x=14 and y=6 16) x=15 and y=5 17) x=16 and y=4 18) x=17 and y=3 19) x=18 and y=2 20) x=19 and y=1 21) x=20 and y=0 By using above value, calculate the x-y = 4 . Then the value which x-y = 4 is the answere. Thus, x=12 and y=8. --------------------------------------… SECOND STEP:- It is the very easy step. Calculate the substraction of the equation x+y = 20 and x-y=4 x + y = 20 - x - y = 4 -------------------------------- y + y = 16 Therefore, y + y = 16 => 2y = 16 => y = 16/2 => y=8 The value of y is calculated. So, lets find the value of x 1) x + y = 20 => x + 8 = 20 => x = 20 - 8 => x = 12 2) x - y = 4 => x - 8 = 4 => x = 4 + 8 => x = 12 Therefore x=12 Thus x =12 and y = 8.

OpenStudy (anonymous):

Add the first equation to the second equation: x+y +x-y -------- 2x=24 Since the ys cancel, we have 2x=24. Dividing both sides by two, we obtain x=12. Plugging in x=12 to the original equations: 12-y=4 →y is 8

OpenStudy (anonymous):

thanks